OpenStudy (anonymous):

Determine the volume of the solid obtained by rotating the portion of the region bounded by and that lies in the first quadrant about the y-axis.

OpenStudy (amistre64):

your trying to tempt me arent ya :)

OpenStudy (anonymous):

Nope. Not in the least. ;)

OpenStudy (amistre64):

is this saying the volume of the first quadrant spun around the y axis?

OpenStudy (anonymous):

Yup. Which is to say, bound by both axis with the limits being each intersection

OpenStudy (amistre64):

i appear to be missing a function to determine proper bound lol; or have gone blind :)

OpenStudy (anonymous):

Oh you're right they didn't post... My bad! It's y=(x^1/3) and y =x/4

OpenStudy (amistre64):

x^3 = x/4 would be the bounds then right...

OpenStudy (amistre64):

cbrt(x)=x/4 typoed it lol

OpenStudy (amistre64):

im going with 0 to 4

OpenStudy (amistre64):

gotta redo that lol

OpenStudy (amistre64):

0 to 8 ...

OpenStudy (anonymous):

Yeah try again lol

OpenStudy (anonymous):

There you go

OpenStudy (amistre64):

my stupidity got in the way

OpenStudy (anonymous):

I forgive you. 64 is awfully close to one. ;)

OpenStudy (amistre64):

spin around the y axis....

OpenStudy (amistre64):

we can do this from 0 to 2 on the y right

OpenStudy (anonymous):

Less bored?

myininaya (myininaya):

lol nerds!

OpenStudy (amistre64):

less.... :)

OpenStudy (anonymous):

Yes, if you mean in terms of y.

OpenStudy (amistre64):

one radius is x=4y the other is .....

OpenStudy (amistre64):

y^3?

OpenStudy (anonymous):

I wouldn't do it that way.

OpenStudy (amistre64):

\[\pi \int\limits_{0}^{2} [4y]^2 - [y^3]^2 dy\]

OpenStudy (anonymous):

Though yes. You can. The top one being?

OpenStudy (amistre64):

why is it gonna break?

OpenStudy (anonymous):

Haha no,but yeah that's right. Let's pretend y just put inthe calculator and got it right because you can push buttons well. :)

OpenStudy (amistre64):

i get 128/3 - 128/7

OpenStudy (amistre64):

512/21 = 24.381 ??

OpenStudy (amistre64):

if I do shelss; i wonder if i get the same result :)

OpenStudy (anonymous):

That was fun :) it's 512 pi/21you forgot pi lol

OpenStudy (amistre64):

i think i dropped a pi lol

OpenStudy (amistre64):

ack!!.... yeah, just realized it :)

OpenStudy (amistre64):

76.59 closer?

OpenStudy (amistre64):

yes!! shell and disk are the same yay!!

OpenStudy (amistre64):

\[2\pi \int\limits_{0}^{8} x(x^{4/3}) -x(\frac{x}{4}) dx\] = 76.59